Grade 9Physical Chemistry

Expression Of Avogadro`s Low By Kinectic Thoary Of Gasses

Kabir
9 Years agoGrade 9

1 Answer

Vikas TU
9 Years ago
Avagadro’s law states that the Volume of the container containing gas is directly proportional to the moles of gas.
v a n
or
v/n = k
And hence,
V1/n1 = V2/n2
This will be valid whenever the Pressure and the Temperature would be constant.